OpenStudy (anonymous):

i need help with my H.W 1. Write the equation of the line below in slope intercept form.

OpenStudy (anonymous):

Slope Intercept: y=mx+b

OpenStudy (anonymous):

@malical THTATS WAT IT LLOOOKS LIKE

OpenStudy (anonymous):

So one point is (0,-2) and another is (-6,0). y=mx+b The b is easy. It is when the lines hits the y-axis. That is at (-2). The m is the slope: m=\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] So: (-2)-(0) / (0)-(-6) -2/6 -1/3 So: y= mx+b y=(-1/3)x-2